OpenStudy (anonymous):

A rectangle is twice as long as it is wide. The perimeter is 114 feet. Find the dimensions.

OpenStudy (anonymous):

Perimeter of a rectangle is 2(length)+2(width) P=2L+2w L=2w P=114 So: 114=2(2w)+2w 114=6w 19=w Now plug into the equation I had for L: L=2w L=2(19) L=38